How to solve 1. 4 x +6 x =9 x ? 2. x-2/x-1 = 1-2/x-1 ?

How to solve   1. 4x+6x=9x ?


                    2. x-2/x-1  = 1-2/x-1  ?

Grade:12

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi Giridhar,

22x+2x3x=32x

divide both sides by 2x3x

(2/3)x+1=(3/2)x

1/y+1=y

y2-y-1=0

y=(1+√5)/2 _ve is not valid.

(3/2)x=(1+√5)/2

x=[log(1+√5)/2]/[log 3/2]